OpenStudy (anonymous):

PLZ HELP THANKS! Find the partial fraction decomposition for the rational expression. 8x+236 OVER (x+7)(x+2)(x+13)

OpenStudy (anonymous):

you want the snap way?

OpenStudy (anonymous):

ha i dont know the best way i ges

OpenStudy (anonymous):

best way is snap way. write it down nice and large. you want to know what number goes over the factor (x+7)

OpenStudy (anonymous):

x + 7 = 0 means x = -7

OpenStudy (anonymous):

put replace x everywhere you see it by -7, but ignore the factor of x + 7 in the denominator

OpenStudy (anonymous):

o ok

OpenStudy (anonymous):

get \[\frac{8\times -7+236}{(-7+2)(-7+13)}\]

OpenStudy (anonymous):

this is what goes over the x + 7 term

OpenStudy (anonymous):

is that the final answer?

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

that is just A here \[\frac{A}{x+7}\]

OpenStudy (anonymous):

so we just found a now we plug that in for?

OpenStudy (anonymous):

yup that is A

OpenStudy (anonymous):

lets make sure it is clear what the goal is. you want to write \[\frac{8x+263}{(x+7)(x+2)(x+13)}=\frac{A}{x+7}+\frac{B}{x+2}+\frac{C}{x+13}\]

OpenStudy (anonymous):

so your job is finding A, B, and C

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so we can find A as i said. put your hand over the x+7 part and just replace x everywhere by -7

OpenStudy (anonymous):

then how i find the others?

OpenStudy (anonymous):

did you get -6?

OpenStudy (anonymous):

or did i mess up?

OpenStudy (anonymous):

o I was not woorking it I was just trying to get how to then was going to do the whole thing

OpenStudy (anonymous):

maybe it is -36 by calculator is off

OpenStudy (anonymous):

so how would i find the others after i find a? haha

OpenStudy (anonymous):

ok lets assume we can do it. the computation is \[\frac{-7\times 8+236}{(-7+2)(-7+13)}\]

OpenStudy (anonymous):

we do the others exactly the same way

OpenStudy (anonymous):

we want to know what goes over x+2

OpenStudy (anonymous):

x+2 = 0 means x = -2. put your hand over that factor and then replace all x's by -2

OpenStudy (anonymous):

o ok!

OpenStudy (anonymous):

the computation is \[\frac{8\times -2+236}{(-2+7)(-2+13)}\]

OpenStudy (anonymous):

and for the third guess what?

OpenStudy (anonymous):

\[\frac{8\times -13+236}{(-13+7)(-13+2)}\]

OpenStudy (anonymous):

first one goes over x+7 second goes over x+2 third goes over x+13

OpenStudy (anonymous):

apparently i am too stupid to actually calculate these numbers but they are the ones you want

OpenStudy (anonymous):

i will try them, let me know what you get

OpenStudy (anonymous):

then we can write the "final answer"

OpenStudy (anonymous):

ha ok ok, let me see if i got this you take one x value from bottom and find x then plug it in to ever x in the problem and do that with all of them then you have three problems and do you solve thos and how?

OpenStudy (anonymous):

then you are done. you need to find A, B and C

OpenStudy (anonymous):

then just write the answer.

OpenStudy (anonymous):

\[A=\frac{8\times (-7)+236}{(-7+2)(-7+13)}\]

OpenStudy (anonymous):

whatever that number is, that is A the number that goes over x + 7

OpenStudy (anonymous):

if i had a better calculator i could compute it quickly.

OpenStudy (anonymous):

how do i do it without calculator or really bad calculator?

OpenStudy (anonymous):

well -8*7=-56

OpenStudy (anonymous):

236-56=180

OpenStudy (anonymous):

numerator is 180

OpenStudy (anonymous):

(-7+2)(-7+13)=-5*-5=25

OpenStudy (anonymous):

and 180/25=36/5

OpenStudy (anonymous):

i guess. did i do it right?

OpenStudy (anonymous):

no i didn't

OpenStudy (anonymous):

(-7+2)(-7+13)=-5*6=-30

OpenStudy (anonymous):

180/-30=-6

OpenStudy (anonymous):

so A = -6 and one term will be \[\frac{-6}{x+7}\]

OpenStudy (anonymous):

\[B=\frac{8\times(-2)+236}{(-2+7)(-2+13)}\]

OpenStudy (anonymous):

what will the last term be over 13?

OpenStudy (anonymous):

the last one will be over x+13

OpenStudy (anonymous):

yes and that is c?

OpenStudy (anonymous):

yes C

OpenStudy (anonymous):

we didn't compute it yet

OpenStudy (anonymous):

did you get B?

OpenStudy (anonymous):

could we

OpenStudy (anonymous):

i wrote it above, do you see it? i just didn't compute it yet

OpenStudy (anonymous):

no this one looks harder then i will do b

OpenStudy (anonymous):

ok let me know what you get and i will check it

OpenStudy (anonymous):

i got a number for C

OpenStudy (anonymous):

ha not yet

OpenStudy (anonymous):

4/x+2

OpenStudy (anonymous):

use a calculator it is easier that way

OpenStudy (anonymous):

oh that is B i didn't compute that one hold on

OpenStudy (anonymous):

i computed C

OpenStudy (anonymous):

ya wht u get

OpenStudy (anonymous):

2

OpenStudy (anonymous):

positive>?

OpenStudy (anonymous):

i did get 4 for B yes

OpenStudy (anonymous):

yes i got positive 2

OpenStudy (anonymous):

ok thanks!

OpenStudy (anonymous):

i may need help later to

OpenStudy (anonymous):

welcome. hope method is clear. they get harder.

OpenStudy (anonymous):

crap haha